OpenStudy (anonymous):

Jesse rents a moving van for $75 & must pay $2 per mile. The following week, Alex rents the same van, is charged $80 for the rental and $1.50 per mile. If they each paid the same amount & drove the same number of miles, how far did they each travel?

OpenStudy (anonymous):

75+2x=80+1.5x 2x-1.5x=80-75 .5x=5 x=5/.5 x=10 miles